OpenStudy (anonymous):

3/4 of a number increased by 12 is equal to5 times a number

OpenStudy (shubhamsrg):

(3/4)*x + 12= 5*x

OpenStudy (anonymous):

how would I find x

OpenStudy (shubhamsrg):

Multiply by 4 on both sides, tell me what you get.

OpenStudy (anonymous):

12n+12=20n

OpenStudy (shubhamsrg):

Nopes, did it wrong, try again/

OpenStudy (shubhamsrg):

(4* ( 3/4) *x) + (4*12) = (4*20*x)

OpenStudy (shubhamsrg):

I mean 4*5*x, on RHS

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

3n+ 48=20n

OpenStudy (shubhamsrg):

yep, seems right. Now subtract 3n from both sides.

OpenStudy (anonymous):

seems right?

OpenStudy (shubhamsrg):

I mean right till now.. SO what do you get after subtracting 3n from both sides?

OpenStudy (anonymous):

48=17n

OpenStudy (shubhamsrg):

correct.

OpenStudy (anonymous):

n=2.82

OpenStudy (shubhamsrg):

bingo ! ^_^

OpenStudy (anonymous):

thanks :)